Find the equation of the line that has the given slope and passes through the given point. Write the equation in the slope-intercept form.
m=-3/4, (3,2)

Answer by Mathtut(3670)

use point slope form y-k=m(x-h) where m is the slope and (h,k) is any point on the line
:
y-2=(-3/4)(x-3)
:
4y-8=-3x+9......muliplied all terms by 4 to eliminate fraction
:
3x+4y=1